A particle is launched such that it's maximum range is 26.4m. what is the speed at which it is launched?


1
Expert's answer
2021-04-07T10:48:53-0400

To be given in question

Rmax=26.4 meter


"2\theta=90°" (maximum range)

To be asked in question

U=?

We know that


"R=\frac{ {u^2}{sin2\theta}}{g}"

"R=\frac{u^2}{g}"


"u^2= Rg"

"u=\sqrt{Rg}"

"g=9.8meter\/sec^2"

Put value

"u=\sqrt{26.4\times9.8}"

"u =16.08" "meter\/sec"